# Tomographic reconstruction of a disease transmission landscape via GPS recorded random paths

### v3

@@ 1.1 Epidemiological background
Predicting spatial disease risk is crucial for managing human, livestock, and wildlife diseases (Morris & Blackburn 2016). Traditionally, spatial prediction of infectious diseases has been understudied (Ostfeld et al. 2005; Kirby et al. 2017), but advances in algorithms and technology have improved spatial inferences of disease dynamics (Ostfeld et al. 2005; Riley et al. 2015; Kirby et al. 2017). Many studies focus on directly transmitted diseases, yet numerous significant epidemics and zoonoses are indirectly transmitted, relying on vectors or environmental reservoirs (Hollingsworth et al. 2015). Examples include vector-borne diseases like malaria and Lyme disease (Torres-Sorando & Rodriguez 1997; Killilea et al. 2008; Messina et al. 2015), and non-vector-borne diseases like anthrax and cholera (Ostfeld et al. 2005).
Inferences about the risk of indirectly transmitted diseases often depend on the spatial distribution of cases or prior knowledge about their ecology and dynamics. Spatially explicit compartmental models require prior knowledge of transmission rates or hypotheses about disease spread, which may not be available for emerging diseases (Harvell et al. 2002; Patz et al. 2008). Understanding how animals use the landscape is essential for managing these diseases (Dinh et al. 2020; Orange, Dinh, Peters, Wisely & Blackburn 2021), as it helps describe the distribution and spatial autocorrelation of animal locations over time (de Thoisy et al. 2020; Albery et al. 2022; Dougherty et al. 2018).
Common methods like Maxent and GARP link climatic conditions to the physiological constraints of pathogens and vectors to predict disease presence (Ahmed et al. 2015; Elith et al. 2011; Stockwell 1999). However, predicting disease risk based solely on presence data is limited, as these points are typically identified infected individuals (Ostfeld et al. 2005). Incorporating animal home range and habitat selection has proven useful in identifying transmission risk of wildlife zoonoses (Ragg & Moller 2000; Dougherty et al. 2022). Coupling individual disease status with animal movements can better identify high-risk areas (Ostfeld et al. 2005), yet host movements are seldom included in disease niche modeling or spatial risk prediction.
Recent approaches have begun using movement data from infected individuals to infer areas of higher infection risk (Dougherty et al. 2018; Brdar et al. 2016; Frias-Martinez et al. 2012). While some methods include spatially explicit compartmental models (Lima et al. 2015), they often rely on underlying assumptions about disease dynamics. In this study, we propose a novel method to map disease risk using animal movement and health status without relying on prior ecological assumptions. We demonstrate this method with simulations and data from native and exotic cervid species infected with Epizootic Hemorrhagic Disease Virus (EHDV) on a Florida hunting ranch.
@@ 1.2 Cervid movement and EHDV prevalence data collection
To investigate disease transmission, we used data from 26 GPS collared WTD (8 females, 18 males) tracked for an average of 149 days during 2015 and 2016. Individuals were captured and GPS collared in the spring, ahead of the EHDV transmission season (Dinh et al., 2020) and recaptured in fall to remove the collars. GPS collars were programmed to collect a GPS location every 15 or 120 minutes (Cauvin et al., 2020; Dinh et al., 2020). Animals were immobilized with cartridge or air powered darts, monitored, collared, bled, and released (see Cauvin et al., 2020; Dinh et al., 2020 for details about capture procedures). A blood sample was obtained from individual at the time of initial capture and again when GPS collars were recovered. Each blood sample was tested for the presence of EHDV-1, EHDV-2 and EHDV-6 antibody titers using a virus neutralization test performed at Texas A&M University (Stallknecht et al., 1996; see Cauvin et al., 2020 for details). We classified animals as positive or negative for each virus type using previous cutoffs (Cauvin et al., 2020). Four individuals were tracked during both years allowing us to use the tracks as independent replicates to increase our sample size to 30 animals (Appendix C). We pre-processed the raw GPS tracks and excluded those of seropositive individuals at both collection times. All tracks were interpolated to 15 minutes. At the end the data consisted of 27 tracks for EHDV-1 infected animals, 22 for EHDV-2, and 30 for EHDV-6. The deer capture and handling protocols were developed by JKB and the ranch wildlife manager and approved by the Institutional Animal Care and Use Committee at the University of Florida (UF IACUC Protocols #201508838 and #201609412).
1.3 Data transformation
Figure 1: An analogy between Positron Emission Tomography (left) and the tomographic reconstruction of a disease propensity map (right): Deer moving in a ranch represent gamma rays traveling through a patient. Whether a deer gets infected is like the signals detected in PET. The goal is to reconstruct the high-propensity area, similar to identifying a tumor.
If, for a second, we imagine that these deer could be trained to enter the ranch at designed locations, walk along lines $L_{i}$ and exits the ranch at the end of their straight walks, then the information whether the deer contracted the disease along these lines could help map the disease propensity on the ranch and each deer walking along a line would essentially have calculated the Radon transform $R_{\mu}(L_{i})=\int_{L_{i}}\mu(u,v)\sqrt{du^{2}+dv^{2}}$ of the propensity map $\mu(u,v)$ as a function of locations $(u,v)$, along the $L_{i}$, for deer and line $i=1,\ldots,n$. This is what positron emission tomography (PET) is doing with rays instead of deer, for non-invasive bran imaging for instance (Figure 1).
While this is not feasible with deer because they do not move along straight lines, pet-imaging would be possible with well trained domesticated animals, but at the price of training them to walk along predetermined straight lines. Deer cannot be trained, but, while moving freely along random paths $P_{i}$, deer gather information about $\mu(u,v)$ with the Radon transform $R_{\mu}(P_{i})=\int_{P_{i}}\mu(u,v)\sqrt{du^{2}+dv^{2}}$, for deer and path $i=1,\ldots,n$. If one can extract information from GPS locations of individuals moving freely in a medium along random paths and contracting a disease or not, then one performs some imaging of the medium, here the disease propensity map.
Going back to the deer movement data, each deer has a blood test at time 00 leading to concentrations of four different antibodies $c_{i,m}(0)$, and at time $t_{i}$ leading to concentrations $c_{i,m}(t_{i})$, $m=1,\ldots,3$ (EHDV-1, EHDV-2 and EHDV-6 antibodies). From these concentrations, the binary responses $y_{i,m}=1_{\{c_{i,m}(t_{i})>c_{i,m}(0)\}}$ declares whether or not the $i$th individual got infected by virus $m$ during the time period, for $i=1,\ldots,n$ and $m=1,\ldots,3$. Individuals move differently in the fenced area $\Omega$, as observed by the GPS tracking system. To map the spatial variability in infection propensity $\mu$, the fenced area $\Omega$ is spatially descretized and partitioned over equal small units $\omega^{j,k}$ indexed by latitude $j$ and longitude $k$, for $(j,k)\in{\cal F}$. The level of discretization $p=|{\cal F}|$ is chosen by the practitioner: the larger $p$, the finer the spatial discretization, but the more parameters to estimate. Correspondingly, each area $\omega^{j,k}$ has the propensity $\mu(\omega^{j,k})$ of contaminating a species with a disease during a unit time period, where $\mu$ is a function of localization $\omega^{j,k}$ in $\Omega$. Using the GPS recordings, one can, for each individual $i$, aggregate the total time $x^{j,k}_{i}$ spent in each small area $\omega^{j,k}$ between its starting time until time $t_{i}$; so let ${\bf x}_{i}=\{x_{i}^{j,k}\}_{(j,k)\in{\cal F}}$ be the vector of total time spent in location $\omega^{j,k}$ for individual $i$, for $i=1,\ldots,n$. The goal is to estimate the propensity maps ${\boldsymbol{\mu}}_{m}=\{\mu_{m}(\omega^{j,k})\}_{(j,k)\in{\cal F}}$ ($m=1,\ldots,3$, one map per EHDV virus) based on the information contained in the data $(y_{i},{\bf x}_{i})_{i=1,\ldots,n}$. We performed this data transformation process (we provide the resultant $X$ and $y$ for each dataset in the supplementary materials) to the EHDV-1, EHDV-2 and EHDV-6 datasets.
1.4 Our approach
Individuals randomly moving in the field are analogous to tomographic rays traveling across a medium, and the resulting binary output indicating whether or not the animals became infected are akin to sensor readings (Figure 1). So we aim at retrieving the disease propensity map, in the spirit of inverse problems such as positron emission tomography (PET). To the best of our knowledge, our approach is novel to disease mapping and indirect risk estimation for the data we described.
@@ 1.4 Our approach
Following on the data preprocessing proposed in Section 1.3, we present our approach in the following way. In Section 2, we model the data with a generalized linear model regularized by a total variation penalty to enforce spatial coherence in the propensity map, we discuss the selection of the regularization parameter and propose an algorithm to solve the convex optimization problem, we derive a statistical test for a constant propensity map, and we show how to perform uncertainty quantification. In Section 3, we perform a Monte Carlo simulation to empirically compare three methods of estimation with three different asymptotics: increasing sample size, increasing spatial resolution, and increasing frequency recording. In Section 4, we analyze the deer data of Section 1.2. In Section 5, we summarize our contributions and point to generalizations of our approach. The research is reproducible (see Supplementary Material).
2 Total variation regularization of GLM
@@ 2.1 Model
Since the three EHDV serotype antibodies considered have their own specificities, three separate models are estimated, so we call the disease marker $y_{i}\in\{0,1\}$ (instead of $y_{i,1}$ for EHDV-1, $y_{i,2}$ for EHDV-2 and $y_{i,3}$ for EHDV-6) for $i=1,\ldots,n$. The classical approach for binary outcomes sees $y_{i}$ as a realization of the random variable $Y_{i}\sim\operatorname{Bernoulli}(\epsilon_{i})$, where the probability $\epsilon_{i}$ of being infected depends on where the $i$th individual spent its time in the fenced area. In particular $\epsilon_{i}$ should be high if the $i$th individual spent long periods of time in areas of high infection propensity; in other words, $\epsilon_{i}$ should be high when high total times $x_{i}^{j,k}$ are observed in regions $\omega^{j,k}$ where the unknown propensity $\mu(\omega^{j,k})$ is high. A simple yet realistic model is the generalized linear model (Nelder & Wedderburn, 1972) which assumes $g(\epsilon_{i})={\bf x}_{i}^{\rm T}{\boldsymbol{\mu}}$, where the logit link $g(\epsilon)=\log(\epsilon/(1-\epsilon))$ maps $(0,1)$ into ${\mathbb{R}}$. Letting ${\bf Y}=(Y_{1},\ldots,Y_{n})$ and $X$ be the $n\times p$ matrix which $i$th row is ${\bf x}_{i}^{\rm T}$, the model is
@@ 2.2 Estimation
The vector of infection propensity ${\boldsymbol{\mu}}$ has length $p=|{\cal F}|$, the cardinality of the lattice system that segments the fenced area $\Omega$. In our application the number of individuals is small, so we choose a fairly coarse segmentation of the fenced area into $p=200$ cells. Regularization is needed, and owing to the spatial structure of the problem, we impose a smoothness constraint on ${\boldsymbol{\mu}}$. Possible approaches models $\mu$ as a linear combination of splines (Wahba, 1990) or wavelets (Donoho, 1995); kernel smoothing is another approach. We instead use total variation (TV) regularization (Rudin et al., 1992) for several reasons: first the domain is not rectangular, second epidemiologists prefer to identify constant propensity regions and third the propensity map may have some peaks (e.g., feeders) and discontinuities (e.g., along a river or a lake). TV provides a solution to these challenges as it easily adapts to a non-regular lattice, it performs segmentation (as it fits a piecewise constant function $\mu$) and it allows sudden peaks and jumps.
To derive TV for our inverse problem, we define a system of neighborhood. An entry $\mu_{l}$ of the vector of infection propensity ${\boldsymbol{\mu}}$ corresponds to a small region $\omega^{j,k}$ in $\Omega$. Each region $(j,k)$ is associated to an entry of ${\boldsymbol{\mu}}$, call it $l$. Each region $(j,k)$ (associated to $l$) has neighbors: call $\partial l$ the set of entries of ${\boldsymbol{\mu}}$ corresponding to the neighbors of $(j,k)$. Because of the spatial structure of ${\boldsymbol{\mu}}$ which maps the propensity of catching a virus somewhere in the ranch, one believes that the value of $\mu_{l}$ is close to $\mu_{l^{\prime}}$ for $l^{\prime}\in\partial l$; here we consider the north-south-east-west neighborhood of a cell. TV imposes smoothness by solving
@@
where $L$ is the likelihood associated to (1), ${\bf y}$ is the vector of infection indicator per individual, $X$ is total time matrix per individual (lines) spent in area $\omega^{j,k}$ (column), and ${\boldsymbol{\mu}}$ is the two-dimensional surface to reconstruct. The regularization parameter $\lambda>0$ controls the smoothness of the estimate $\hat{\boldsymbol{\mu}}_{\lambda}$ solution to (2). One standard selection of $\lambda$ is cross-validation, which requires $n$ large for a stable estimation. In our situation, the number $n$ of individuals is rather small, so we employ the quantile universal threshold (Giacobino et al., 2017) that is geared towards estimating and testing the propensity map ${\boldsymbol{\mu}}$. The quantile universal threshold is based on calibrating a choice of $\lambda$ under the null hypothesis $H_{0}$ that the propensity map is constant (that is, no regions are more infectious than others): its goal is to retrieve a constant map with high probability under $H_{0}$. This choice of $\lambda$ remains performant under alternative hypotheses as supported by the LASSO theory (Tibshirani, 1996; Bühlmann & van de Geer, 2011) and total variation (Rudin et al., 1992; Sardy & Monajemi, 2019). Deriving the quantile universal threshold for our TV-estimator (2) is based on the following property.
Property 2.1.
@@
where $h({\boldsymbol{\mu}};{\bf y})=-\log{\rm L}(g^{-1}(X{\boldsymbol{\mu}});{\bf y})$ is the negative log-likelihood associated to model (1).
@@
Since $\alpha$ corresponds to the false discovery rate (of detecting a non-constant map) under $H_{0}$, one chooses $\alpha$ small, say $\alpha=0.05$. The distribution of $\Lambda$ is unknown however, so $\lambda^{\rm QUT}_{\alpha}=F^{-1}_{\Lambda}(1-\alpha)$ has no closed form expression. Instead $\lambda^{\rm QUT}_{\alpha}$ can be estimated by Monte Carlo, simulating $m$ realizations ${\bf y}^{(1)},\ldots,{\bf y}^{(m)}$ of ${\bf Y}\sim\operatorname{Bernoulli}(g^{-1}(\beta_{0}X{\bf 1}))$ for some $\beta_{0}$, for instance $\hat{\beta}_{0}=\operatorname*{\arg\!\min}_{\beta_{0}\in{\mathbb{R}}}h(\beta_{0}{\bf 1};{\bf y})$ for the data ${\bf y}$ at hand. Then by calculating the corresponding $\lambda_{0}({\bf y}^{(1)},X),\ldots,\lambda_{0}({\bf y}^{(m)},X)$ and by taking their empirical upper $\alpha$-quantile, one estimates $\lambda^{\rm QUT}_{\alpha}$. The zero-thresholding function implicitly defined in Property 2.1 requires solving the non-trivial optimization problem in (3); fortunately, (3) can be rewritten as an easily solvable linear program, namely

### v4

@@ 1.1 Epidemiological background
Predicting spatial disease risk is critical for managing infectious diseases that affect humans, livestock, and wildlife (Albery et al., 2022; Miller et al., 2013; Morris & Blackburn, 2016). Although this area has historically been underexplored, recent advances in computational methods, animal tracking technologies, and modeling approaches have significantly improved our ability to infer the spatial dynamics of disease transmission (Dougherty et al., 2018; White et al., 2018; Manlove et al., 2022; Dougherty et al., 2022). Traditionally, research has focused on directly transmitted diseases, despite the fact that many major outbreaks and emerging zoonoses—such as anthrax, brucellosis, and vector-borne infections—are transmitted indirectly through environmental reservoirs or vectors (Altizer et al., 2011; Rayl et al., 2021; Dougherty et al., 2022). Spatial predictions for such indirectly transmitted diseases often rely on aggregated case data or prior ecological knowledge. However, these approaches may be inadequate for emerging or rapidly evolving disease systems (Altizer et al., 2011; Manlove et al., 2022).
Mechanistic and spatially explicit compartmental models offer more formal approaches, but typically require system-specific information, such as transmission rates or assumptions about host-environment interactions, which may be unavailable for novel pathogens (Dougherty et al., 2018; Roosa & Chowell, 2019; Glennon et al., 2021; Dankwa et al., 2022). Consequently, understanding how wildlife moves across landscapes and interacts with environmental reservoirs or vectors has become increasingly important for identifying areas of elevated disease risk (Manlove et al., 2022).
Modern species distribution modeling methods—such as Maxent, GARP, and boosted regression trees—have been used to associate disease risk with environmental or climatic covariates (Elith et al., 2011; Ahmed et al., 2015; Solano-Villarreal et al., 2019; Dougherty et al., 2022). However, these presence-only models rarely account for host movement dynamics and behavioral heterogeneity (Dougherty et al., 2018, 2022). Incorporating home range analyses and GPS-based movement data has proven effective for mapping zoonotic disease hotspots and clarifying transmission routes (Rayl et al., 2021). Linking infection status with high-resolution movement data—such as that derived from telemetry—enables researchers to identify high-risk spatial zones and critical time windows for pathogen transmission (Dougherty et al., 2018; Wilber et al., 2022; Malmberg et al., 2025). Nevertheless, this approach remains underutilized in disease niche modeling.
Recent studies have increasingly integrated animal movement data to identify areas of elevated infection risk and support targeted surveillance and intervention strategies (Dekelaita et al., 2023; Ciss et al., 2023; McDuie et al., 2024). While spatially explicit compartmental models have improved risk prediction (White et al., 2018), these approaches typically require a priori knowledge of disease ecology or model-specific parameters (Manlove et al., 2022; Gao et al., 2023; Akter & Deardon, 2025) . Other models rely on detailed information about interactions between infected and susceptible individuals (Wilber et al., 2022). While these frameworks represent major advances, they still depend on assumptions about disease transmission pathways and host-environment relationships.
In this study, we present and validate a novel method for mapping infectious disease risk that combines individual-level movement trajectories and health status data without requiring prior knowledge of the transmission process or environmental context. This method can serve as a powerful stand-alone tool or be integrated with environmental datasets to link spatial risk predictions with ecological mechanisms of disease transmission (Manlove et al., 2022). We demonstrate the utility of this approach using both simulated data and a case study of native and exotic cervids infected with epizootic hemorrhagic disease virus (EHDV) on a Florida hunting ranch.
@@ 1.2 Cervid movement and EHDV prevalence data collection
To investigate disease transmission, we used data from 26 GPS collared WTD (8 females, 18 males) tracked for an average of 149 days during 2015 and 2016. Individuals were captured and GPS collared in the spring, ahead of the EHDV transmission season (Dinh et al., 2020) and recaptured in fall to remove the collars. GPS collars were programmed to collect a GPS location every 15 or 120 minutes (Cauvin et al., 2020; Dinh et al., 2020). Animals were immobilized with cartridge or air powered darts, monitored, collared, bled, and released (see Cauvin et al., 2020; Dinh et al., 2020 for details about capture procedures). A blood sample was obtained from individual at the time of initial capture and again when GPS collars were recovered. Each blood sample was tested for the presence of EHDV-1, EHDV-2 and EHDV-6 antibody titers using a virus neutralization test performed at Texas A&M University (Stallknecht et al., 1996; see Cauvin et al., 2020 for details). We classified animals as positive or negative for each virus type using previous cutoffs (Cauvin et al., 2020). Four individuals were tracked during both years allowing us to use the tracks as independent replicates to increase our sample size to 30 animals (See Supplementary materials). We pre-processed the raw GPS tracks and excluded those of seropositive individuals at both collection times. All tracks were interpolated to 15 minutes. At the end the data consisted of 27 tracks for EHDV-1 infected animals, 22 for EHDV-2, and 30 for EHDV-6. The deer capture and handling protocols were developed by JKB and the ranch wildlife manager and approved by the Institutional Animal Care and Use Committee at the University of Florida (UF IACUC Protocols #201508838 and #201609412).
1.3 Our contributions
We propose a novel framework to estimate the landscape propensity of EHDV infection using only deer movement data, without relying on external environmental or ecological covariates. Individual deer wandering across the landscape are conceptualized as tomographic rays traversing a medium, while the corresponding binary infection outcomes serve as sensor readings (Figure 1). We regularize the high dimensional inverse problem with a total-variation approach and derive the mathematical and practical guidelines for choosing the penalty parameter, yielding models that remain readily interpretable for practitioners.
@@ 1.3 Our contributions
In Section 2 we describe a data-transformation procedure, inspired by the Radon transform in positron-emission tomography (PET), that recasts the task as an epidemiological tomographic inverse problem. Because the number of GPS-collared deer is typically small, this transformation produces a high-dimensional inference problem with limited observations. In Section 3, we model the data with a generalized linear model regularized by a total variation penalty to enforce spatial coherence in the propensity map. In this front, we derive the necessary mathematical result for an efficient selection of the regularization parameter with the quantile universal threshold, that does not require cross-validation. We propose an algorithm to solve the convex optimization problem, we derive a statistical test for a constant propensity map, and we show how to perform uncertainty quantification. In Section 4, we explore alternative tomographic approaches, and we also introduce a potential subsequent data aggregation designed to accommodate direct image smoothing methods. In Section 5, we perform a Monte Carlo simulation to empirically compare all methods with three different asymptotics: increasing sample size, increasing spatial resolution, and increasing frequency recording. In Section 6, we analyze the deer data of Section 1.2. In Section 7, we summarize our contributions and point to generalizations of our approach. The research is reproducible (see Supplementary Materials).
2 Tomographic data transformation
Figure 1: An analogy between Positron Emission Tomography (left) and the tomographic reconstruction of a disease propensity map (right): Deer moving in a ranch represent gamma rays traveling through a patient. Whether or not a deer gets infected is like a positron is emitted in PET. The goal is to reconstruct the high-propensity area, similar to identifying a tumor.
If, for a second, we imagine that the deer described in Section 1.2 could be trained to enter the ranch at designed locations, walk along lines $L_{i}$ and exits the ranch at the end of their straight walks, then the information whether the deer contracted the disease along these lines could help map the disease propensity on the ranch and each deer walking along a line would essentially have calculated the Radon transform $R_{\mu}(L_{i})=\int_{L_{i}}\mu(u,v)\sqrt{du^{2}+dv^{2}}$ of the propensity map $\mu(u,v)$ as a function of locations $(u,v)$, along the $L_{i}$, for deer and line $i=1,\ldots,n$. This is what positron emission tomography (PET) is doing with rays instead of deer, for non-invasive brain imaging for instance (Figure 1).
This is not feasible with deer because they do not move along straight lines but, while moving freely along random paths $P_{i}$, deer gather information about $\mu(u,v)$ with the Radon transform $R_{\mu}(P_{i})=\int_{P_{i}}\mu(u,v)\sqrt{du^{2}+dv^{2}}$, for deer and path $i=1,\ldots,n$. If one can extract information from GPS locations of individuals moving freely in a medium along random paths and contracting a disease or not, then one performs some imaging of the medium, here the disease propensity map.
Going back to the deer movement data, each deer has a blood test at time 0 leading to concentrations of four different antibodies $c_{i,m}(0)$, and at time $t_{i}$ leading to concentrations $c_{i,m}(t_{i})$, $m=1,\ldots,3$ (EHDV-1, EHDV-2 and EHDV-6 antibodies). From these concentrations, the binary responses $y_{i,m}=1_{\{c_{i,m}(t_{i})>c_{i,m}(0)\}}$ declares whether or not the $i$th individual got infected by virus $m$ during the time period, for $i=1,\ldots,n$ and $m=1,\ldots,3$. Individuals move differently in the fenced area $\Omega$, as observed by the GPS tracking system. To map the spatial variability in infection propensity $\mu$, the fenced area $\Omega$ is spatially descretized and partitioned over equal small units $\omega^{j,k}$ indexed by latitude $j$ and longitude $k$, for $(j,k)\in{\cal F}$. The level of discretization $p=|{\cal F}|$ is chosen by the practitioner: the larger $p$, the finer the spatial discretization, but the more parameters to estimate. Correspondingly, each area $\omega^{j,k}$ has the propensity $\mu(\omega^{j,k})$ of contaminating a species with a disease during a unit time period, where $\mu$ is a function of localization $\omega^{j,k}$ in $\Omega$. Using the GPS recordings, one can, for each individual $i$, aggregate the total time $x^{j,k}_{i}$ spent in each small area $\omega^{j,k}$ between its starting time until time $t_{i}$; so let ${\bf x}_{i}=\{x_{i}^{j,k}\}_{(j,k)\in{\cal F}}$ be the vector of total time spent in location $\omega^{j,k}$ for individual $i$, for $i=1,\ldots,n$. The goal is to estimate the propensity maps ${\boldsymbol{\mu}}_{m}=\{\mu_{m}(\omega^{j,k})\}_{(j,k)\in{\cal F}}$ ($m=1,\ldots,3$, one map per EHDV virus) based on the information contained in the data $(y_{i},{\bf x}_{i})_{i=1,\ldots,n}$. We performed this data transformation process (we provide the resultant $X$ and $y$ for each dataset in the Supplementary materials) to the EHDV-1, EHDV-2 and EHDV-6 data. Individuals randomly moving in the field resemble tomographic rays passing through a medium, with their binary infection outcomes acting like sensor measurements (Figure 1). So we aim at retrieving the disease propensity map, in the spirit of inverse problems such as positron emission tomography (PET). To the best of our knowledge, our approach is novel to disease mapping and indirect risk estimation for the data we described.
3 Regularization of GLM with total variation and the quantile universal threshold
@@ 3.1 Model
Since the three EHDV serotype antibodies considered have their own specificities, three separate models are estimated, so we call the disease marker $y_{i}\in\{0,1\}$ for $i=1,\ldots,n$ (instead of $y_{i,1}$ for EHDV-1, $y_{i,2}$ for EHDV-2 and $y_{i,3}$ for EHDV-6). The classical approach for binary outcomes sees $y_{i}$ as a realization of the random variable $Y_{i}\sim\operatorname{Bernoulli}(\epsilon_{i})$, where the probability $\epsilon_{i}$ of being infected depends on where the $i$th individual spent its time in the fenced area. In particular $\epsilon_{i}$ should be high if the $i$th individual spent long periods of time in areas of high infection propensity; in other words, $\epsilon_{i}$ should be high when high total times $x_{i}^{j,k}$ are observed in regions $\omega^{j,k}$ where the propensity $\mu(\omega^{j,k})$ is high. A simple yet realistic model is the generalized linear model (Nelder & Wedderburn, 1972) which assumes $g(\epsilon_{i})={\bf x}_{i}^{\rm T}{\boldsymbol{\mu}}$, where the logit link $g(\epsilon)=\log(\epsilon/(1-\epsilon))$ maps $(0,1)$ into $\mathbb{R}$}. Letting ${\bf Y}=(Y_{1},\ldots,Y_{n})$ and $X$ be the $n\times p$ matrix which $i$th row is ${\bf x}_{i}^{\rm T}$, the model is
@@ 3.2 Estimation
The vector of infection propensity $\boldsymbol{\mu}$} has length $p=|{\cal F}|$, the cardinality of the lattice system that segments the fenced area $\Omega$. In our application the number of individuals is small, so we choose a fairly coarse segmentation into $p=200$ cells. Regularization is needed, and owing to the spatial structure of the problem, smoothness can be imposed on $\boldsymbol{\mu}$} in many ways (see Section 4). We choose total variation (TV) smoothing (Rudin et al., 1992) for several reasons: first the domain is not rectangular, second epidemiologists prefer to identify piecewise-constant propensity regions and third the propensity map may have some peaks (e.g., feeders) and discontinuities (e.g., along a river or a lake). TV provides a solution to these challenges as it easily adapts to a non-regular lattice, it performs segmentation (as it fits a piecewise constant function) and it allows sudden peaks and jumps. We also propose a novel simple rule to select the smoothing parameter $\lambda$.
TV requires a system of neighborhood in the propensity map $\boldsymbol{\mu}$}. An entry $\mu_{l}$ of the vector of infection propensity $\boldsymbol{\mu}$} corresponds to a small region $\omega^{j,k}$ in $\Omega$. Each region $(j,k)$ is associated to an entry of $\boldsymbol{\mu}$}, call it $l$. Each region $(j,k)$ (associated to $l$) has neighbors: call $\partial l$ the set of entries of $\boldsymbol{\mu}$} corresponding to the neighbors of $(j,k)$. Because of the spatial structure of $\boldsymbol{\mu}$} which maps the propensity of catching a virus somewhere in the ranch, one believes that the value of $\mu_{l}$ is close to $\mu_{l^{\prime}}$ for $l^{\prime}\in\partial l$; here we consider the north-south-east-west neighborhood of a cell. TV imposes smoothness by solving
@@
where $L$ is the likelihood associated to (1), $\bf y$} is the vector of infection indicator per individual, $X$ is total time matrix per individual (lines) spent in area $\omega^{j,k}$ (column), and $\boldsymbol{\mu}$} is the two-dimensional surface to reconstruct. The regularization parameter $\lambda>0$ controls the smoothness of the estimate $\hat{\boldsymbol{\mu}}_{\lambda}$ solution to (2). One standard selection of $\lambda$ is cross-validation, which requires $n$ large for a stable estimation. In our situation, the number $n$ of individuals is rather small, so we employ the quantile universal threshold (Giacobino et al., 2017) that is geared towards estimating $\boldsymbol{\mu}$} rather than good prediction of $X{\boldsymbol{\mu}}$. The quantile universal threshold is based on calibrating a choice of $\lambda$ under the null hypothesis $H_{0}$ that the propensity map is constant (that is, no regions are more infectious than others): its goal is to retrieve a constant map with high probability under $H_{0}$. This choice of $\lambda$ remains performant under alternative hypotheses as supported by the LASSO theory (Tibshirani, 1996; Bühlmann & van de Geer, 2011) and total variation (Rudin et al., 1992; Sardy & Monajemi, 2019), and demonstrates strong empirical performance in practical applications (Diaz-Rodriguez et al., 2021). Deriving the quantile universal threshold for our tomographic TV-estimator (2) requires the following property.
Property 3.1.
@@
where $h({\boldsymbol{\mu}};{\bf y})=-\log{\rm L}(g^{-1}(X{\boldsymbol{\mu}});{\bf y})$ is the negative log-likelihood associated to (1), and $D$ is the $q\times p$ matrix of the $q$ finite differences such that $\|D{\boldsymbol{\mu}}\|_{1}=\sum_{l=1}^{p}\sum_{l^{\prime}\in\partial l}|\mu_{l}-\mu_{l^{\prime}}|$.
@@
Since $\alpha$ corresponds to the false discovery rate (of detecting a non-constant map) under $H_{0}$, one chooses $\alpha$ small, say $\alpha=0.05$. The distribution of $\Lambda$ is unknown however, so $\lambda^{\rm QUT}_{\alpha}=F^{-1}_{\Lambda}(1-\alpha)$ has no closed form expression. Instead $\lambda^{\rm QUT}_{\alpha}$ can be estimated by Monte Carlo, simulating $m$ realizations ${\bf y}^{(1)},\ldots,{\bf y}^{(m)}$ of ${\bf Y}\sim\operatorname{Bernoulli}(g^{-1}(\beta_{0}X{\bf 1}))$ for some $\beta_{0}$, for instance $\hat{\beta}_{0}=\operatorname*{\arg\!\min}_{\beta_{0}\in{\mathbb{R}}}h(\beta_{0}{\bf 1};{\bf y})$ for the data $\bf y$} at hand. Then by calculating the corresponding $\lambda_{0}({\bf y}^{(1)},X),\ldots,\lambda_{0}({\bf y}^{(m)},X)$ and by taking their empirical upper $\alpha$-quantile, one estimates $\lambda^{\rm QUT}_{\alpha}$. The zero-thresholding function implicitly defined in Property 3.1 requires solving the non-trivial optimization problem in (3); fortunately, (3) can be rewritten as an easily solvable linear program, namely